# Diffusion Models Are Innate One-Step Generators

### v1

@@ Abstract
Diffusion Models (DMs) have achieved great success in image generation and other fields. By fine sampling through the trajectory defined by the SDE/ODE solver based on a well-trained score model, DMs can generate remarkable high-quality results. However, this precise sampling often requires multiple steps and is computationally demanding. To address this problem, instance-based distillation methods have been proposed to distill a one-step generator from a DM by having a simpler student model mimic a more complex teacher model. Yet, our research reveals an inherent limitations in these methods: the teacher model, with more steps and more parameters, occupies different local minima compared to the student model, leading to suboptimal performance when the student model attempts to replicate the teacher. To avoid this problem, we introduce a novel distributional distillation method, which uses an exclusive distributional loss. This method exceeds state-of-the-art (SOTA) results while requiring significantly fewer training images. Additionally, we show that DMs’ layers are activated differently at different time steps, leading to an inherent capability to generate images in a single step. Freezing most of the convolutional layers in a DM during distributional distillation leads to further performance improvements. Our method achieves the SOTA results on CIFAR-10 (FID 1.54), AFHQv2 64x64 (FID 1.23), FFHQ 64x64 (FID 0.85) and ImageNet 64x64 (FID 1.16) with great efficiency. Most of those results are obtained with only 5 million training images within 6 hours on 8 A100 GPUs. This breakthrough not only enhances the understanding of efficient image generation models but also offers a scalable framework for advancing the state of the art in various applications.
@@ 1 Introduction
Diffusion Models (DMs, Song & Ermon (2020a),Song & Ermon (2020b),Ho et al. (2020)) have led to an explosion of generative vision models. By sampling through the trajectory defined by the Probability Flow Ordinary Differential Equation (PF ODE) or Stochastic Differential Equation (SDE) solvers (Song et al. (2021)), DMs can generate high-quality images and videos, surpassing many traditional methods, including Generative Adversarial Network (GAN, Goodfellow et al. (2014)). However, to produce high-quality results, DMs require multiple sample steps to generate an accurate trajectory, which is computationally expensive. To alleviate this problem, a series of distillation methods (Salimans & Ho (2022), Song et al. (2023), Kim et al. (2024)) have been proposed to distill a one-step generative model. However, these methods often require substantial computational resources, yet produce poorer performance than original models.
In this work, we first demonstrate that the underlying limitation of these instance-based distillation methods is due to the fact that the local minima for the teacher model and the student model can be very different. We then introduce a distributional distillation method based on GAN loss. Our method is highly efficient, reaching state-of-the-art (SOTA) results with only 5 million training images. In comparison, instance-based distillation methods often require hundreds of millions of training images.
@@ 3.1 GAN Distillation at Distribution level
For distilling a one-step generator from DMs, we introduce a new method called GDD (GAN Distillation at Distribution Level) , which uses only a distribution-level loss WITHOUT any instance-level supervision. Importantly, unlike in SDXL-Lightning or DMD, the discriminator in our method receives real data instead of the data generated by the teacher model as the ground truth. We use non-saturating GAN objective defined as:
@@ 4.2 Achieving One-Step Generation via Freezing
The results confirm our idea (Table 2). Freezing most of the convolutional layers in the UNet blocks not just works. It actually produces better performance. Further freezing the QKV projections (Vaswani et al. (2023)) causes a slight decrease in the performance, and freezing the skip layers on the residual connections leads to worse results, suggesting that tuning of these two types of layers is necessary for distillation. These results strongly support our theory that the ability for one-step generation is inherent in DMs: with most of the original parameters locked (85.8% of the parameters contained in the convolutional layers) in the pre-trained score model, our model achieves results better than the previous SOTA results. We term the method that freezes most of the convolutional layers as GDD-I (GAN Distillation at the Distribution level using Innate ability).
@@ 4.4 Comprehensive Performance
Finally, we perform a comprehensive test of both GDD and GDD-I on CIFAR-10, AFHQv2 64x64 (Choi et al. (2020)), FFHQ 64x64 (Karras et al. (2019)), and ImageNet 64x64 (Deng et al. (2009)). The results for class conditioning generation are reported on CIFAR-10 and ImageNet 64x64. The pre-trained diffusion models are from EDM. We use the ADM (Dhariwal & Nichol (2021)) architecture for ImageNet and the NCSN++ architecture for the other datasets. We use lr=1e-4 for CIFAR-10, lr=2e-5 for AFHQv2 64x64 and FFHQ 64x64, and lr=8e-6 for ImageNet. Batch size is set to 512 for ImageNet and 256 for all other datasets. More details can be found in Appendix B.4. We report FID for all datasets, Inception Score (IS, Salimans et al. (2016)) for CIFAR-10, precision and recall (Kynkäänniemi et al. (2019)) for ImageNet 64x64. For the discriminator, we use VGG16 with batch normalization and EfficientNet-lite0 as the feature network for CIFAR-10, DeiT (Touvron et al. (2021)) and EfficientNet-lite0 for the other datasets. The optimal settings in Section 3.2 is used.
@@
Table 4: Comprehensive comparison between GDD, GDD-I, and previous work. Tests are done on ImageNet 64x64, with class conditioning.
@@ Results.
The results can be found in Table 3 and 4. Both GDD and GDD-I achieve the SOTA results on all datasets that we have tested. Our methods are not only suitable for unconditional distillation, but also show great performance on class conditioning generation tasks. Moreover, GDD-I outperforms GDD by a significant margin in most of the datasets, except AFHQv2 64x64. On CIFAR-10, our methods achieve the best FID of 1.54 and the best IS of 10.11. Similarly, superior performance is found on AFHQv2 64x64 (FID=1.23) and FFHQ 64x64 (FID=0.85). In ImageNet 64x64, the model also achieves SOTA performance (FID=1.16, precision=0.77). Together, these results demonstrate the robustness and effectiveness of our methods. They support our hypothesis that DMs are capable of generating results in one-step innately.
@@ 6 Conclusion
In this work, we have identified an inherent limitation in instance-based distillation methods, stemming from the disparities between the teacher and the student models. To circumvent this limitation, we first propose a novel method, GDD (GAN Distillation in Distribution level), which leverages an exclusive distributional loss to efficiently distill a one-step generator from diffusion models. We further investigate the activation patterns of diffusion model layers, and develop GDD-I (GDD using Innate ability), which enhances performance by freezing most of the convolutional layers in the GDD model during the distillation process and outperforms the original GDD method. Our work not only establishes new SOTA results with minimal computational resources but also elucidates a critical principle in the field of diffusion distillation, paving the way for more efficient and effective model training methodologies in image generation and beyond.
@@ D LR 1e-4 4e-5 4e-5 4e-5
We distill EDMs for all datasets and methods for fair comparison. EMA decay is applied to the weights of generator for sampling following previous work. EMA decay rate is calculated by EMA Halflife and EMA warmup is used as in EDM. EMA warmup ratio is set to 0.05 for all datasets, and this leads to a gradually grow EMA decay rate. We use Adam optimizer with $\beta_{1}=0,\beta_{2}=0.99$ without weight decay for both the generator and the discriminator. No gradient clip is applied. Mixed precision training is adopted for all experiments with BFloat16 data type, and results with Float16 and Float32 are similar. We apply no learning rate scheduler. Images are resized to 224 first before they are fed to the PG discriminator or CD-LPIPS loss. All models are trained on a cluster of Nvidia A100 GPUs. Hyperparameters used for GDD and GDD-I training can be found in Table 6. GDD and GDD-I share the same hyperparameters except for the number of training images, since GDD-I will not overfit when GDD gets worse FID.

### v2

@@ Abstract
Diffusion Models (DMs) have achieved great success in image generation and other fields. By fine sampling through the trajectory defined by the SDE/ODE solver based on a well-trained score model, DMs can generate remarkable high-quality results. However, this precise sampling often requires multiple steps and is computationally demanding. To address this problem, instance-based distillation methods have been proposed to distill a one-step generator from a DM by having a simpler student model mimic a more complex teacher model. Yet, our research reveals an inherent limitations in these methods: the teacher model, with more steps and more parameters, occupies different local minima compared to the student model, leading to suboptimal performance when the student model attempts to replicate the teacher. To avoid this problem, we introduce a novel distributional distillation method, which uses an exclusive distributional loss. This method exceeds state-of-the-art (SOTA) results while requiring significantly fewer training images. Additionally, we show that DMs’ layers are differentially activated at different time steps, leading to an inherent capability to generate images in a single step. Freezing most of the convolutional layers in a DM during distributional distillation enables this innate capability and leads to further performance improvements. Our method achieves the SOTA results on CIFAR-10 (FID 1.54), AFHQv2 64x64 (FID 1.23), FFHQ 64x64 (FID 0.85) and ImageNet 64x64 (FID 1.16) with great efficiency. Most of those results are obtained with only 5 million training images within 6 hours on 8 A100 GPUs.
@@ 1 Introduction
Diffusion Models (DMs, Song & Ermon (2020a),Song & Ermon (2020b),Ho et al. (2020)) have led to an explosion of generative vision models. By sampling through the trajectory defined by Probability Flow Ordinary Differential Equation (PF ODE) or Stochastic Differential Equation (SDE) solvers (Song et al. (2021)), DMs can generate high-quality images and videos, surpassing many traditional methods, including Generative Adversarial Networks (GAN, Goodfellow et al. (2014)). However, to produce high-quality results, DMs require multiple sample steps to generate an accurate trajectory, which is computationally expensive. To alleviate this problem, a series of distillation methods (Salimans & Ho (2022), Song et al. (2023), Kim et al. (2024)) have been proposed to distill a one-step generative model. However, these methods often require substantial computational resources, yet produce poorer performance than original models.
In this work, we first demonstrate that the underlying limitation of these instance-based distillation methods is due to the fact that the local minima for the teacher model and the student model can be very different. We then introduce a distributional distillation method based exclusively on GAN loss. Our method is highly efficient, reaching state-of-the-art (SOTA) results with only 5 million training images. In comparison, instance-based distillation methods often require hundreds of millions of training images.
@@ 3.1 GAN Distillation at Distribution level
For distilling a one-step generator from DMs, we introduce a new method called GDD (GAN Distillation at Distribution Level) , which uses only a distribution-level loss WITHOUT any instance-level supervision. Importantly, unlike in SDXL-Lightning or DMD, our method uses real data instead of the data generated by the teacher model as the ground truth. We use non-saturating GAN objective defined as:
@@ 4.2 Achieving One-Step Generation via Freezing
The results confirm our idea (Table 2). Freezing most of the convolutional layers in the UNet blocks not just works. It actually produces better performance. Further freezing the QKV projections (Vaswani et al. (2023)) causes a slight decrease in the performance, and freezing the skip layers on the residual connections leads to worse results, suggesting that tuning of these two types of layers is necessary for distillation. These results strongly support our theory that the ability for one-step generation is inherent in DMs: with most of the original parameters locked (85.8% of the parameters contained in the convolutional layers) in the pre-trained score model, our model achieves results better than the previous SOTA results. We term the method that freezes most of the convolutional layers as GDD-I (GAN Distillation at the Distribution level using Innate ability). One of the critical challenges in training GANs is the inherent instability of the training process. However, our method circumvents this issue by freezing the majority of parameters in both the Discriminator and the Generator. As a result, the training process of GDD-I remains remarkably stable, with minimal instances of mode collapse.
@@ 4.4 Comprehensive Performance
Finally, we perform a comprehensive test of both GDD and GDD-I on CIFAR-10, AFHQv2 64x64 (Choi et al. (2020)), FFHQ 64x64 (Karras et al. (2019)), and ImageNet 64x64 (Deng et al. (2009)). The results for class-conditional generation are reported on CIFAR-10 and ImageNet 64x64. The pre-trained diffusion models are from EDM. We use the ADM (Dhariwal & Nichol (2021)) architecture for ImageNet and the NCSN++ architecture for the other datasets. We use lr=1e-4 for CIFAR-10, lr=2e-5 for AFHQv2 64x64 and FFHQ 64x64, and lr=8e-6 for ImageNet. Batch size is set to 512 for ImageNet and 256 for all other datasets. More details can be found in Appendix B.4. We report FID for all datasets, Inception Score (IS, Salimans et al. (2016)) for CIFAR-10, precision and recall (Kynkäänniemi et al. (2019)) for ImageNet 64x64. For the discriminator, we use VGG16 with batch normalization and EfficientNet-lite0 as the feature network for CIFAR-10, DeiT (Touvron et al. (2021)) and EfficientNet-lite0 for the other datasets. The optimal settings in Section 3.2 is used.
@@
Table 4: Comprehensive comparison between GDD, GDD-I, and previous work. Tests are done on ImageNet 64x64, class-conditional.
@@ Results.
The results can be found in Table 3 and 4. Both GDD and GDD-I achieve the SOTA results on all datasets that we have tested. Our methods are not only suitable for unconditional distillation, but also show great performance on class-conditional generation tasks. Moreover, GDD-I outperforms GDD by a significant margin in most of the datasets, except AFHQv2 64x64. On CIFAR-10, our methods achieve the best FID of 1.54 and the best IS of 10.11. Similarly, superior performance is found on AFHQv2 64x64 (FID=1.23) and FFHQ 64x64 (FID=0.85). In ImageNet 64x64, the model also achieves SOTA performance (FID=1.16, precision=0.77). Together, these results demonstrate the robustness and effectiveness of our methods. They support our hypothesis that DMs are capable of generating results in one-step innately.
@@ 6 Conclusion
In this work, we have identified an inherent limitation in instance-based distillation methods, stemming from the disparities between the teacher and the student models. To circumvent this limitation, we first propose a novel method, GDD (GAN Distillation in Distribution level), which leverages an exclusive distributional loss to efficiently distill a one-step generator from diffusion models. We further investigate the activation pattern of diffusion model layers and develop GDD-I (GDD using Innate ability), which enhances performance by freezing most of the convolutional layers in the GDD model during the distillation process and outperforms the original GDD method. Our work not only establishes new SOTA results with minimal computational resources but also elucidates a critical principle in the field of diffusion distillation, paving the way for more efficient and effective model training methodologies in image generation and beyond.
@@ D LR 1e-4 4e-5 4e-5 4e-5
We distill EDMs for all datasets and methods for fair comparison. EMA decay is applied to the weights of generator for sampling following previous work. EMA decay rate is calculated by EMA Halflife and EMA warmup is used as in EDM. EMA warmup ratio is set to 0.05 for all datasets, and this leads to a gradually grow EMA decay rate. We use Adam optimizer with $\beta_{1}=0,\beta_{2}=0.99$ without weight decay for both the generator and the discriminator. No gradient clip is applied. For class-conditional generation, we use no classifier-guidance, but using simple class embedding in discriminator following StyleGAN2-XL. Mixed precision training is adopted for all experiments with BFloat16 data type, and results with Float16 and Float32 are similar. We apply no learning rate scheduler. Images are resized to 224 first before they are fed to the PG discriminator or CD-LPIPS loss. All models are trained on a cluster of Nvidia A100 GPUs. Hyperparameters used for GDD and GDD-I training can be found in Table 6. GDD and GDD-I share the same hyperparameters except for the number of training images, since GDD-I will not overfit when GDD gets worse FID.